OpenStudy (roberts.spurs19):

Please help! Find the position vector 'a' of a point on the line described by the equation that follows \[\frac{ x+1 }{ 3} = \frac{ y+4 }{-3 }=\frac{ z }{ 2 }\]

OpenStudy (anonymous):

You can express all the (x,y,z) coordinates in function of only one, (x+1)/3 = (y+4)/-3, make y in function of x (x+1)/3 = z/2, make z in function of x Now that you have these functions, create the vectorial function that describes the position in space of all the points in this line in function of coordinate x

OpenStudy (roberts.spurs19):

thank you!

OpenStudy (anonymous):

Np ;)